OpenStudy (anonymous):

Find work done by F=<√(lnx)−y,x^2+e^(tany)> around the triangle from (0,0) to (1,1), to (-1,1), and back to (0,0). How should I begin to do this beast? It seems like I should use Green's thm, but how?

OpenStudy (anonymous):

According to green's thm, it seems I need to find Nx-My, which I found to be 2x+1 So, \[\int\limits_{}^{}\int\limits_{}^{} (2x+1)da\] but if the initial and end points are (0,0) then that makes no sense.

OpenStudy (anonymous):

I originally had the \[\int\limits_{0}^{1}\int\limits_{-1}^{1}2x+1 dxdy\] and my professor wrote "your H.S. region is a rectangle." Would that help?